### v1

# Emergent morphogenesis via planar fabrication enabled by a reduced model of composites

Abstract
The ability to engineer complex three-dimensional shapes from planar sheets with precise, programmable control underpins emerging technologies in soft robotics, reconfigurable devices, and functional materials. Here, we present a reduced-order numerical and experimental framework for a bilayer system consisting of a stimuli-responsive thermoplastic sheet (Shrinky Dink) bonded to a kirigami-patterned, inert plastic layer. Upon uniform heating, the active layer contracts while the patterned layer constrains in-plane stretch but allows out-of-plane bending, yielding programmable 3D morphologies from simple planar precursors. Our approach enables efficient computational design and scalable manufacturing of 3D forms with a single-layer reduced model that captures the coupled mechanics of stretching and bending. Unlike traditional bilayer modeling, our framework collapses the multilayer composite into a single layer of nodes and elements, reducing the degrees of freedom and enabling simulation on a 2D geometry. This is achieved by introducing a novel energy formulation that captures the coupling between in-plane stretch mismatch and out-of-plane bending—extending beyond simple isotropic linear elastic models. Experimentally, we establish a fully planar, repeatable fabrication protocol using a stimuli-responsive thermoplastic and a laser-cut inert plastic layer. The programmed strain mismatch drives an array of 3D morphologies, such as bowls, canoes, and flower petals, all verified by both simulation and physical prototypes.
pacs:
Valid PACS appear here
I Introduction
Rapid and cost-effective manufacturing of 3D structures from planar 2D materials is challenging and the manufacturing methods are in significant demand in fields such as soft robotics, automobile manufacturing, biomedical devices, aerospace components, and deployable structures. The existing structures fabricated from active and responsive materials have enabled applications in aerospace furuya1992concept ; hanaor2001evaluation ; santiago2013advances , medicine du2018programmable ; banerjee2020origami ; babaee2019temperature , and intelligent machines wang2017modular ; chi2022bistable , but they remain primarily material-focused. There is a growing need for scalable manufacturing frameworks that can accelerate real-world deployment. In this context, computational design – particularly approaches that enable rapid simulation and optimization – plays a pivotal role. Our work addresses this challenge by introducing a reduced computational model that enables efficient, accurate simulation of shape-morphing composites, leveraging algorithms originally developed in the computer graphics community grinspun2003discrete .
A variety of pathways have been explored to drive shape morphing, including tailored material designs dudek2025shape , composite structures daynes2013review ; rodriguez2016shape , microstructurally engineered polymers aharoni2018universal ; jin2022shape ; studart2014bioinspired , origami-inspired mechanisms hawkes2010programmable ; tolley2014self , and architected metastructures boston2022spanwise ; risso2022highly ; zhang2025shape . The transformation from 2D to 3D forms is governed by the interplay between stretching and bending armon2011geometry , where strain mismatches embedded in planar preforms can lead to a variety of 3D shapes aharoni2018universal ; ware2015voxelated ; giudici2022multiple ; modes2011gaussian . In particular, bilayer composites comprising a stimuli-responsive material bonded to a patterned kirigami layer mungekar2024directed offer a scalable,
manufacturing-driven strategy for programmable shape morphing across a broad range of applications, such as deployable structures, patient-specific biomedical devices, and adaptive surfaces. In this work, we use a bilayer system composed of a Shrinky Dink thermoplastic and an inert kirigami-patterned plastic as a model platform for programmable morphable surfaces. Our reduced-order numerical method, demonstrated on this system, can be readily augmented to simulate and guide the design of other classes of shape-morphing composite shells.
Figure 1: Examples demonstrating the validation between 3D shapes manufactured by physical experiments using the protocol developed here and the ones generated by numerical simulations using the proposed energy in Equation 6. Scalebar: 1 cm.
In this paper, we introduce a numerical framework for efficient and accurate prediction of stimulus-driven 3D shape morphing in bilayer composite shells. Unlike conventional approaches that require explicit modeling of multiple layers, we represent the entire layered composite using a single layer of nodes and formulate an effective energy model that captures the coupled mechanics of the bilayer system. Our method builds on the discrete shells approach grinspun2003discrete , where the total elastic energy of a thin structure is computed from in-plane stretching (changes in edge lengths) and out-of-plane bending (dihedral angles between adjacent triangular elements), each referenced to rest lengths and rest angles that define the undeformed, stress-free configuration. We advance this framework by allowing these rest lengths and rest angles to evolve explicitly as functions of the applied stimulus (heat in this work), enabling the model to capture permanent, stimulus-driven deformations such as those caused by thermal contraction in heat-shrinking layers. This energy-based approach yields accurate predictions of complex 3D configurations arising from 2D patterned precursors, without resorting to computationally expensive full 3D simulations. We validate our framework by comparing numerical predictions to physical experiments on bilayer composites composed of a heat-shrinking thermoplastic (Shrinky Dink) and a layer of 3D printed PLA (polylactic acid).
Figure 1 shows representative 3D shapes fabricated in physical experiments (left) and their numerically simulated counterparts (right). The experimental fabrication protocol builds directly on prior work mungekar2024directed , while the present study provides, for the first time, a predictive numerical framework for these systems. Our main contributions are:
  • Introduction of a reduced-order modeling approach in which the composite is represented by a single layer of nodes; this framework is broadly applicable to a variety of layered or architected material systems, beyond bilayer composites.
  • Explicit incorporation of external stimuli by formulating the rest lengths and rest angles (i.e., the stress-free, undeformed configuration) as functions of the applied stimulus, enabling accurate modeling of permanent, stimulus-driven morphing.
  • Quantitative demonstration that the numerical predictions closely match experimental measurements across a range of complex shapes and kirigami patterns; see Figure 1.
We begin by detailing the experimental approach, including the materials, methods, and fabrication protocols. Next, we introduce our new energy formulation and describe its numerical implementation within the discrete shells algorithm. To demonstrate the effectiveness of our approach, we study three representative kirigami patterns and analyze their resulting 3D shapes through both numerical simulations and physical experiments. We quantify the differences between the simulated and experimental shapes at various stages of the stimulus, and find excellent agreement, thereby validating the predictive capability of our framework.
II Results and Discussion
II.1 Emergent 3D Shapes via Planar Fabrication of Bilayer Composites
The experimental approach exploits the rapid contraction of thermally responsive polystyrene (Shrinky Dink) together with the geometric programmability enabled by 3D printing. Polystyrene is selected for its reliable shrinkage (up to 60% reduction in length), low cost, and safe processing in standard laboratory ovens. The shrinkage ratio can be precisely tuned by adjusting the heating time, providing an additional degree of control over the final shape. For programmable morphing, we digitally design and directly 3D print the inert plastic (PLA) layer in the desired kirigami pattern – a network of openings and solid regions that enables localized bending and selectively constrains shrinkage. This additive manufacturing approach ensures high precision and reproducibility, which are difficult to achieve with manual cutting methods such as scissors. The patterned PLA is then bonded to the polystyrene sheet, allowing for spatial control of deformation and a customizable mechanical response.
Figure 2 presents three representative examples demonstrating the morphing of 3D shapes from planar bilayer composites via thermal actuation. In the first row, panels A(i)–C(i) show physical scans of the digitally designed inert kirigami layer (solid black regions) alongside the shapes of the Shrinky Dink substrate (dashed boundaries) for each 2D precursor, while panels A(ii)-C(ii) display the corresponding numerical meshes. The second row shows the resulting physical and numerically simulated 3D shapes. Depending on the kirigami pattern, the flat 2D precursor undergoes stimulus-driven transformation in the oven to form a variety of 3D shapes, such as a bowl, a canoe, and flower petals. The close agreement between the experimentally scanned shapes and the simulation results demonstrates the predictive capability of our numerical framework.
Figure 2: Examples demonstrating the morphing of 2D bilayer composites into 3D shapes via thermal actuation. The first row shows the 2D designs: each design consists of a specific kirigami pattern (PLA layer) on a thermoplastic (Shrinky Dink) substrate of defined shape. The second row displays the resulting 3D shapes. Subfigures A, B, and C correspond to three representative cases: bowl, canoe, and flower petal morphologies, respectively. For each design, the left column (i) presents results from physical experiments, and the right column (ii) shows the corresponding numerical simulations using the new energy model in Equation 6.
To fabricate the morphable bilayer composites, we first design a kirigami pattern for the PLA layer based on heuristics that relate the pattern geometry to the intended 3D morphed shape. The chosen pattern is then digitally modeled and 3D printed as a thin PLA structure. This PLA layer is bonded to the polystyrene (Shrinky Dink) sheet using a two-part Gorilla epoxy. The high bending and stretching stiffness of PLA constrains shrinkage in the bonded regions, allowing the patterned geometry to locally direct where in-plane contraction is allowed or resisted. As a result, the composite exhibits spatially varying stiffness, which can be further tuned by adjusting the thickness of the PLA layer to achieve complex 3D surfaces.
Figure 3 illustrates the complete fabrication process. First, the kirigami pattern is designed based on the desired final shape; for example, Figure 3A shows a six-arm star on a circular thermoplastic sheet. The pattern is then digitally modeled with the prescribed thickness using Fusion fusion2025 and fabricated by 3D printing, as shown in Figure 3B. Next, the printed PLA pattern is bonded to a precisely cut thermoplastic substrate, and the assembled bilayer composite is cured for over 10 hours to ensure robust adhesion (Figure 3C). After curing, thermal actuation is used to induce shape morphing, as shown in Figure 3D, by heating the planar composite in an oven for approximately 15 seconds. The response of the Shrinky Dink layer is characterized by measuring its shrinkage as a function of normalized temperature (Figure 3E), where the temperature is scaled by the glass transition temperature ($T_{g}$) of polystyrene, an amorphous thermoplastic without a sharp melting point. The sheet begins to shrink when $T/T_{g}>1$ (temperatures are measured in Kelvins), but the shrink ratio $L/L_{0}$ does not stabilize until it exceeds 1.15, corresponding to temperatures above $300^{\circ}$F, as indicated by a vertical dashed line in Figure 3E. Consequently, all experiments are performed at temperatures greater than $300^{\circ}$F.
Using this protocol, we explored a range of additively manufactured patterns with varying dimensions and sheet shapes. This paper highlights three representative designs (Figure 1A–C), each defined by a distinct combination of substrate domain and kirigami pattern: (A) a six-arm star kirigami pattern on a circular Shrinky Dink substrate, (B) a two-star kirigami pattern on a rectangular substrate, and (C) a cross kirigami pattern on a square substrate. After thermal activation at $300^{\circ}$F for approximately 15 seconds, the planar composites reliably morph into 3D shapes, which are then reconstructed and analyzed through 3D scanning. In the following section, we present our numerical method for predicting these experimentally observed 3D shapes.
Figure 3: Experimental setup. (A) Design of the kirigami pattern (inert PLA layer) and selection of the substrate domain (thermoplastic Shrinky Dink sheet). (B) 3D printing of the kirigami pattern. (C) Assembly of the bilayer composite by bonding the 3D-printed PLA pattern to the substrate. (D) Thermal actuation in an oven for shape morphing. (E) Measured variation of the shrink ratio ($L/L_{0}$) as a function of normalized temperature ($T/T_{g}$), where $T_{g}=200^{\circ}$F (366.5366.5K) is the glass transition temperature of polystyrene. The sheet begins to shrink when $T/T_{g}>1$, and the shrink ratio stabilizes above $T/T_{g}\approx 1.15$ (indicated by the vertical dashed line). Scale bars in (B), (C), and (D) denote 1 cm.
II.2 Numerical Simulation with a Reduced Model
Motivated by the experimental observations of programmable shape morphing in bilayer composites – comprising a 3D printed PLA kirigami layer bonded to a Shrinky Dink sheet – we propose a new energy formulation that couples planar stretching with out-of-plane deformation. We first briefly review the discrete elastic shells method in Section II.2.1, and then introduce our reduced energy model for composite shells in Section II.2.2.
II.2.1 Discrete Elastic Shells method
The discrete elastic shells framework provides a discrete differential geometry-based approach for modeling thin shells, where the thickness is much smaller than the in-plane dimensions. In this formulation, the structure is represented as a triangular mesh with $N$ nodes connected by $N_{e}$ edges, as shown in Figure 4. The specific meshes used to simulate the three representative examples in this work are shown in Figures 4A–C. Each node (e.g., $a,b,c,d$ in Figure 4D) has three spatial degrees of freedom, corresponding to its position in 3D space, and the full configuration is described by the vector $\mathbf{X}$ of size $3N$. An edge is the connection between two nodes, labeled as 0, 11, 22, 33, and $4$ in Figure 4D. Edges that are shared by two triangles are referred to as hinges (for example, edge 0 is the common edge of triangles abcabc and adbadb). The mesh contains $N_{h}$ such hinges. At every hinge, the angle between the normals of the adjacent triangles is defined as the dihedral angle. Edges can undergo stretching, and hinges can undergo bending; these deformations store stretching and bending energies, respectively. The corresponding elastic energy expressions are introduced later in this section.
The equations of motion for the DOF vector $\mathbf{X}$ are given by
$\mathbb{M}\ddot{\mathbf{X}}+\nabla E=\mathbf{F}_{\textrm{ext}},$ (1)
where $\mathbb{M}$ is the (diagonal) mass matrix, $\dot{\mathbf{X}}$ and $\ddot{\mathbf{X}}$ denote the first and second time derivatives of $\mathbf{X}$, $\nabla E$ is the gradient of the total elastic energy with respect to the DOF vector (i.e., the negative of the elastic force), and $\mathbf{F}_{\textrm{ext}}$ is the external force vector. To advance the simulation from $t=t_{k}$ to $t_{k+1}=t_{k}+\Delta t$, we employ the implicit Euler method and discretize Equation 1 as
$\mathbb{M}\frac{1}{\Delta t}\left(\frac{\mathbf{X}(t_{k+1})-\mathbf{X}(t_{k})}{\Delta t}-\dot{\mathbf{X}}(t_{k})\right)+\nabla E(\mathbf{X}(t_{k+1}))=\mathbf{F}_{\textrm{ext}}(t_{k+1}).$ (2)
Given the positions $\mathbf{X}(t_{k})$ and velocities $\dot{\mathbf{X}}(t_{k})$ at the previous time step, the new positions $\mathbf{X}(t_{k+1})$ are computed by solving this nonlinear system via the Newton–Raphson method. The Jacobian matrix for each Newton iteration is
$\mathbf{J}=\frac{\mathbb{M}}{\Delta t^{2}}+\nabla^{2}E(\mathbf{X})-\nabla\mathbf{F}_{\textrm{ext}},$ (3)
where $\nabla^{2}E$ is the Hessian of the elastic energy with respect to nodal positions. Since our study focuses on the static equilibrium shape, the inertial term $\mathbb{M}\ddot{\mathbf{X}}$ is neglected and the external force $\mathbf{F}_{\textrm{ext}}(t_{k+1})$ is set to zero. In this case, the final configuration corresponds to a minimum of the total elastic energy. The velocity at $t_{k+1}$ can be calculated simply from $\dot{\mathbf{X}}(t_{k+1})=\left(\mathbf{X}(t_{k+1})-\mathbf{X}(t_{k})\right)/\Delta t$.
In the discrete shells framework, the total elastic energy is the sum of stretching energies (associated with each edge) and bending energies (associated with each hinge) savin_growth_2011 :
$E\bigl{(}\bm{X}\bigr{)}=\sum_{i=1}^{N_{e}}\frac{1}{2}\,k_{s}\,\varepsilon_{i}^{2}+\sum_{j=1}^{N_{h}}\frac{1}{2}\,k_{b}\,\left(\theta_{j}-\bar{\theta}_{j}\right)^{2},$ (4)
where $\varepsilon_{i}$ (axial strain for edge $i$) and $\theta_{j}$ (dihedral angle at hinge $j$) are defined below as macroscopic, geometrically invariant strain measures, and $\bar{\theta}_{j}$ denotes the rest dihedral angle at hinge $j$, i.e., the dihedral angle in the reference (undeformed) configuration. For a flat plate, $\bar{\theta}_{j}=0$, while for a curved shell, $\bar{\theta}_{j}$ may be nonzero and spatially varying.
The strains are fully determined by the DOF vector $\mathbf{X}$ and are invariant with respect to the global coordinate system. Specifically, the axial strain for each edge is $\varepsilon_{i}=l_{i}/l_{0,i}-1$, where $l_{i}$ is the current edge length and $l_{0,i}$ is its rest (reference) length in the undeformed mesh. The dihedral angle $\theta_{j}$ is defined as the angle between the normals of the two triangles sharing hinge $j$. The stretching and bending stiffnesses are
$k_{s}$ $=\frac{\sqrt{3}}{2}\,Y\,h\,l_{0}^{2},$ (5)
$k_{b}$ $=\frac{2}{\sqrt{3}}\,D,\quad D=\frac{Y\,h^{3}}{12}\,.$
where $Y$ is the Young’s modulus, $h$ is the shell thickness, and $l_{0}$ is the mean edge length. These discrete expressions are constructed such that, in the limit of mesh refinement, the model recovers the continuum Föppl–von Kármán equations for nonlinear thin plates. The three composites have domains $\Omega$ of 100 mm radius (pattern A), 100 mm $\times$ 156 mm (pattern B), and 100 mm $\times$ 100 mm (pattern C), Figures 2 and 4. With the mesh configurations in Figure 4, we have $l_{0}=3.2$ $m$, $4.0$ mm, and $5.8$ mm for A, B, and C, respectively. The thicknesses of two layers and their corresponding effective stretching and bending stiffness values, $k_{s}$ and $k_{b}$, are given in Section II.3.
In our model, the computational mesh consists of two distinct types of regions, Figure 4 A-C(i): (i) bilayer regions $\mathcal{B}$, where both the Shrinky Dink and the inert kirigami layer are present, and (ii) single-layer regions $\Omega\setminus\mathcal{B}$, where only the Shrinky Dink substrate remains (corresponding to the cut-out areas of the kirigami pattern). The material properties – such as Young’s modulus $Y$, and consequently the stretching and bending stiffnesses $k_{s}$ and $k_{b}$ – are assigned locally to each mesh element based on its region: bilayer regions use effective properties of the combined layers, while single-layer regions use properties of the Shrinky Dink alone. The specific numerical values used for all material parameters are provided in the Section II.3.
While the assignment of material properties enables accurate representation of spatial heterogeneity in the composite, classical discrete plate models are limited to flat, stress-free reference states. In contrast, discrete shells extend this framework by permitting nonzero natural (rest) curvature, achieved by introducing nonzero rest dihedral angles. This is implemented by replacing the bending term $\frac{1}{2}\,k_{b}\,\theta_{j}^{2}$ with $\frac{1}{2}\,k_{b}\,\left(\theta_{j}-\theta_{j}^{\mathrm{rest}}\right)^{2}$ in the energy expression. Such an extension is essential for capturing the permanent, stimulus-induced shape changes observed in our bilayer composites. Building on these principles, we next present our reduced energy model for programmable composite shells, which incorporates stimulus-dependent rest configurations to predict the full range of morphable shapes observed in experiments.
Figure 4: Reference-configurations meshes for designs A-C in Figures 1 and 2. A-C(i) Thermoplastic PLA bilayer boundaries (solid lines) on discrete mesh. A-C(ii) The distribution of thermal strains as defined in Equation 7. See Section II.2.2 for definitions of $\Omega$, $\mathcal{B}$, $\varepsilon_{pre}^{\mathrm{th}}$, and $d_{i}$. D. Local node and edge IDs for a bending pair. The number of nodes $N_{n}$, number of edges $N_{e}$, and number of triangular elements $N_{t}$, $(N_{n},N_{e},N_{t})$ are (970, 2800, 1831) for pattern A, (1215, 3505, 2291) for pattern B, and (388, 1087, 700) for pattern C, respectively.
II.2.2 Energy Formulation for Stimulus-Responsive Composite Shells
The core innovation of our approach is to extend the discrete shells framework so that the elastic energy depends explicitly on both the DOF vector $\mathbf{X}$ – representing the shape of the structure – and the applied external stimulus, such as temperature. This enables us to efficiently model the permanent, stimulus-driven morphing of bilayer shells using only a single layer of mesh nodes, without the need for computationally intensive 3D elements. The key physical mechanism is the non-uniform thermal strain induced in the bilayer composite: when a 3D-printed PLA pattern is bonded to a Shrinky Dink substrate, only the Shrinky Dink contracts under heating, while the PLA remains largely inert. This mismatch in planar contraction leads to locally varying natural curvatures – captured in our model as stimulus-dependent rest dihedral angles at the mesh hinges. By properly designing the kirigami pattern, we control which regions bend and by how much, enabling programmable shape transformation. Building on the framework established above, we generalize the energy expression in Eq. 4 by allowing the rest dihedral angle $\bar{\theta}_{j}$ to depend explicitly on the applied thermal stimulus, $T$.
In our reduced model, the total elastic energy is written as
$E(\bm{X},T)=\sum_{i=1}^{N_{e}}\frac{1}{2}k_{s}\left[\varepsilon_{i}(\bm{X})-\varepsilon_{i}^{\mathrm{th}}(T)\right]^{2}+\sum_{j=1}^{N_{h}}\frac{1}{2}k_{b}\left[\theta_{j}(\bm{X})-\beta l_{0}\Delta\varepsilon_{j}(\bm{X},T)\right]^{2},$ (6)
where $\varepsilon_{i}^{\mathrm{th}}(T)$ is the stimulus-induced thermal strain in the Shrinky Dink layer (measured from experiments, cf. Figure 3E). The first term accounts for in-plane stretching, while the second term models out-of-plane bending. Here, $\theta_{j}(\bm{X})$ is the dihedral angle at hinge $j$, and $\beta l_{0}\Delta\varepsilon_{j}(\bm{X},T)$ replaces the rest dihedral angle $\bar{\theta}_{j}$ in Eq. 4, quantifying the local mismatch in planar strain across the two triangles sharing each hinge. The parameter $\beta$ controls the strength of this coupling and is determined by fitting to experimental data (see Supporting Information for details). Specifically, $\beta$ (unit $m^{-1}$) denotes the coupling effect per unit triangle‐side length, and $\beta l_{0}$ gives the coupling strength for a mesh whose triangles have a mean side length $l_{0}$. For convenience, we set $\beta=1/l_{0}$ here. With this definition, we aim to diminish the dependence of $\beta$ on numerical discretization. The explicit expression for $\Delta\varepsilon_{j}$ is provided later in this section. We write $\varepsilon_{i}(\bm{X})$, $\theta_{j}(\bm{X})$, and $\Delta\varepsilon_{j}(\bm{X},T)$ to emphasize that these quantities are computed from the DOF vector $\bm{X}$ and, where relevant, the applied stimulus $T$. Similarly, $\varepsilon_{i}^{\mathrm{th}}(T)$ denotes the thermal strain as a function of $T$, computed directly from the experimentally measured contraction of the Shrinky Dink sheet at each temperature (see Figure 3E).
The thermal strain results from a mismatch caused by shrinkage due to temperature change, is described by $\varepsilon_{pre}^{\mathrm{th}}=\alpha\Delta T$. By observing experimental results, we found that deformation around the boundary of the Shrinky Dink sheet is more significant than in the region occupied by the PLA composite. Another contributing factor is the glue effect. In some areas, the glue has spread beyond the PLA layer, making the composite stiffer than the pure Shrinky Dink sheet. Consequently, the magnitude of the thermal shrinkage strain increases with the distance from the PLA region, where it remains zero. This variation of thermal strain is given by
$\begin{gathered}\varepsilon_{i}^{\mathrm{th}}=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{d_{i}}{d_{\max}},\quad d_{i}=\min_{j\in\mathcal{B}}\bigl{\lVert}\mathbf{x}_{i}-\mathbf{x}_{j}\bigr{\rVert},\quad d_{\max}=\max_{i\in\Omega\setminus\mathcal{B}}d_{i},\end{gathered}$ (7)
where $\Omega$ is the set of all edge midpoints in the composite domain. $\mathcal{B}$ is the subset of $\Omega$ lying inside the PLA and Shrinky Dink bilayer region. $\mathbf{x}_{i}$ is the midpoint coordinate of edge $i\in\Omega\setminus\mathcal{B}$, and $\mathbf{x}_{j}$ of edge $j\in\mathcal{B}$. $\varepsilon_{pre}^{\mathrm{th}}$ is the thermal strain that develops in a pure Shrinky Dink sheet. $d_{i}$ is the shortest distance from midpoint $i$ to the PLA and Shrinky Dink bilayer region. $d_{\max}$ is the largest distances $d_{i}$ over the region.
The planar strain mismatch at each hinge is computed as
$\Delta\varepsilon_{j}(\bm{X},T)=\sum_{p=1}^{4}s_{p}[\varepsilon_{jp}(\bm{X},T)-\varepsilon^{th}_{jp}(\bm{X},T)],$ (8)
where $\varepsilon_{jp}$ and $\varepsilon^{th}_{jp}$ are the axial strains and axial thermal strains, respectively, of the four edges adjacent to hinge $j$ indicated by $1,2,3,$ and $4$ in Figure 4D, and $s_{p}\in\{+1,-1\}$ are orientation factors determined by the mesh geometry (see Figure 4D for edge numbering and orientation conventions). Physically, $\Delta\varepsilon_{j}$ quantifies the difference in in-plane contraction between the two triangles sharing hinge $j$, which is the local geometric origin of natural curvature. The corresponding thermal strains, $\varepsilon_{jp}^{\mathrm{th}}(T)$, are computed from the experimentally measured contraction of the Shrinky Dink sheet at each temperature, as reported in Figure 3E.
This reduced energy formulation incorporates both the spatial heterogeneity of the composite and the direct influence of external stimulus. By minimizing $E(\bm{X},T)$ for a given $T$, we predict the equilibrium 3D shape for any kirigami design and actuation protocol. The model is validated in the next section by comparison with 3D scans of the fabricated structures.
II.3 Numerical Results Validated with Experiments
The bilayer composite is modeled as a single-layer elastic shell that incorporates thermal strains and dihedral angles as reference strains via the energy, Equation 6. Each region is characterized by its Young’s modulus $Y$. At room temperature, for the Shrinky Dink sheet, denoted as layer 1, $Y_{1}$ typically ranges from 1GPa to 3GPa mishra_effect_2005 . For the 3D-printed PLA layer, denoted as layer 2, $Y_{2}$ ranges from 2GPa to 4GPa crupano_investigating_2024 . However, at a temperature over $300^{\circ}F$ which is the requirement of the thermal actuation, the modulus becomes around two to three orders of magnitude smaller than their values at room temperature lendvai2023valorization ; tabi2019application ; ricarte2024tutorial . So we set $Y_{1}=1$MPa for the Shrinky Dink, and $Y_{2}=3$MPa for the PLA layer. Given the scaling of stiffness and energy are both linear in Young’s modulus, the unit from GPa to MPa does not affect the solution to equilibrium. The resulting effective stretching stiffness is ugural2011advanced
$k_{s}=\sum_{i=1}^{2}k_{s}^{i},$ (9)
where $k_{s}^{i}$ for each layer is given by Equation 5.
The relation between flexural rigidity and bending stiffness of a plate is $k_{b}=(2/\sqrt{3})D$. For a homogeneous plate, $D=Yh^{3}/12$. For a bilayer composite plate, we have the effective flexural rigidity
$D/b=D_{\rm eff}/b=\sum_{i=1}^{2}Y_{i}\biggl{[}\frac{h_{i}^{3}}{12}+h_{i}(y_{i}-\bar{y})^{2}\biggr{]}.$ (10)
Set the bottom of Shrinky Dink (layer 1) at $y=0$, we can define the following for each layer
$y_{1}=\frac{h_{1}}{2},\quad y_{2}=h_{1}+\frac{h_{2}}{2},\quad A_{1}=bh_{1},\quad A_{2}=bh_{2},$
and the location of the bending neutral axis is ugural2011advanced
$\bar{y}=\frac{Y_{1}A_{1}y_{1}+Y_{2}A_{2}y_{2}}{Y_{1}A_{1}+Y_{2}A_{2}}.$
Recall that layer 1 in Equation 10 denotes the Shrinky Dink sheet, and layer 2 denotes the PLA layer. In our experiments, we used thicknesses of $h_{1}=0.3$mm for the Shrinky Dink sheet, and $h_{2}=0.7$ $m$, $0.6$ mm, and 1.01.0mm for the PLA layer in patterns A, B, and C in Figure 2, respectively. Since stretching is observed more difficult than bending in the experiments, we scaled the Young’s modulus for stretching, and thus $k_{s}$, up by one order of magnitude.
We model the deformation as a static process, solving the equilibrium equations via the Newton-Raphson method. Thermal actuation is applied in incremental load steps with an adaptive step size chosen to satisfy the convergence criterion $|\Delta\mathbb{X}|/|\mathbb{X}|<10^{-5}$. Thermal strains are prescribed according to Equation 7, see Section II.2.2 for justification. Because the Shrinky Dink layer contracts on heating, the induced thermal strains are negative in sign, $\epsilon^{th}_{pre}<0$. To account of practical oven heating, we consider a small, spatially uniform thermal-flux load during the heating process, which decays to zero by the end of the actuation.
We discretize the composite domain into triangular elements, three meshes in the reference configuration are shown in Fig. 4. We have verified the finite-difference approximations of the gradient and Hessian of the total energy, and the analytical expressions for both the gradient and Hessian of the energy are given in the Appendix. The numerical implementation was further verified with analytical solutions for uniaxial tension and compression of a simple geometry domain, and the deflection of a cantilever beam under an endpoint load. Complex behaviors, such as bulging and necking also emerge from simple deformation modes zhang_iterated_2024 . Given these verifications, all subsequent simulations are obtained with the same numerical implementations.
Figure 5: Representative snapshots, starting with the pattern of Figure 2A during the thermally actuated manufacturing process. From the top to the bottom rows, the duration of the thermal stimulus increases. In each row, (i) shows the experimentally scanned shape, (ii) shows the numerical simulation, (iii) shows the distribution of the absolute value of the dihedral angle $|\theta|$, and (iv) shows the distribution of the axial strain $\epsilon$.
Figure 5 presents the experimental scans, numerical simulations, and contours of dihedral angle $\theta$ and axial strains $\epsilon$ for a circular composite with a six-arm PLA pattern, Figure 2A. The term “Stimulus” refers to the monotonically increasing stimulus, which is heat here, during the process. From top to bottom, the simulations are in equilibrium with prescribed thermal strains $\epsilon^{th}_{pre}=-0.20,-0.30$ and 0.77-0.77, respectively. As noted above, these thermal strains are nonuniform and vary with each material point’s distance to the PLA pattern. Larger thermally induced shrinkage results in more significant deformations and axial strains. A greater mismatch in axial strains leads to larger dihedral angles. To accommodate these deformations, the structure morphs from a planar configuration into a 3D shape. Deformation localizes heavily between the arms of the 3D-printed PLA pattern, where the Shrinky Dink layer undergoes the largest shrinkage, indicated by A-C(iii) and A-C(iv) in Figure 5. Overall, excellent agreement is observed between the scanned experimental shapes (first column) and the numerical simulations at each stage of the thermal actuation.
Figure 6: Representative snapshots starting with the pattern of Figure 2B during the thermal actuated manufacturing process. Description of the subfigures is available in the caption of Figure 5.
Figure 6 shows the results for the rectangular composite with a two-star PLA pattern in Figure 2B, subjected to thermal strains $\epsilon^{th}_{pre}=-0.08,-0.15$ and 0.30-0.30, from the top to the bottom, respectively. As the prescribed thermal axial strain increases over time, the structure develops larger axial strains $\epsilon$, strains mismatch, dihedral angles $\theta$, and overall curvature. Deformation localizes near the four corners of the Shrinky Dink layer. The slight asymmetry in the deformed shapes arises from the mesh used in the numerical simulation and minor anisotropies in the bilayer composite due to physical manufacturing defects. Nevertheless, the numerical simulations closely approximate the scanned experimental results.
Figure 7: Representative snapshots starting with the pattern of Figure 2C during the thermal actuated manufacturing process. Description of the subfigures is available in the caption of Figure 5.
Similarly, Figure 7 illustrates the deformed configurations at thermally actuated strains $\epsilon^{th}_{pre}=-0.10$, 0.30-0.30 and 0.50-0.50 for the cross pattern in Figure 2C. As the prescribed thermal strain increases, deformation concentrates increasingly between pattern’s arms. Because the Shrinky Dink layer is square rather than circular, these concentration zones form four sharp “petals” instead of the rounded bowl seen in Figure 5.
These three examples demonstrate excellent agreement between scanned experimental 3D shapes and the numerical simulations morphed from 2D planar composites across various domain geometries, patterns, and levels of thermal actuation. More complex structures can be achieved by assembling multiple shapes and using multiscale modeling. The substantial computational demand of multiscale modeling across multiple time and length scales can be mitigated through iterated learning of operators for out-of-distribution data zhang_iterated_2024 . Potential applications of the three examples include deployable flower petals, bowls, frame-based pillows, canoes, and tent-type structures, each achieved by selecting an appropriate reference length scale.
II.4 Structural similarity index measure
To compare the scanned experimental shapes with numerically simulated ones quantitatively, we use the structural similarity index measure (SSIM), defined at each voxel in a 3D domain. For completeness, we briefly describe our application. We discretize the 3D domain into an $N_{s}\times N_{s}\times N_{s}$ grid and calculate SSIM at each voxel as nilsson2020
$\text{SSIM}=\frac{\bigl{(}2\,\mu_{1}\,\mu_{2}+c_{1}\bigr{)}\,\bigl{(}2\,\sigma_{12}+c_{2}\bigr{)}}{\bigl{(}\mu_{1}^{2}+\mu_{2}^{2}+c_{1}\bigr{)}\,\bigl{(}\sigma_{1}^{2}+\sigma_{2}^{2}+c_{2}\bigr{)}},$ (11)
where $N_{s}=10$. The voxel intensities are normalized to $0,L_{s}]$, and we use the default parameter values $k_{1}=0.01$, $k_{2}=0.03$, $L_{s}=1$, and $c_{i}=(k_{i}*L_{s})^{2}$ for $i=1,2$ nilsson2020 . Here, $\mu_{1}$, $\mu_{2}$ are local means, $\sigma_{1}^{2}$, $\sigma_{2}^{2}$ the variances, and $\sigma_{12}$ the covariance. To make the scanned and simulated shapes comparable, we uniformly scale, translate, and rotate the simulated shapes to align with the scanned experimental data. For our results, the SSIM ranges from 0 to 1, with 0 indicating no structural correlation and 1 indicating perfect structural similarity.
Figure 8 summarizes the SSIM between experimental scans and numerical simulations for all three composite shapes and their morphing processes. Panels (i) in A-C show a side view of each shapes alongside definition of the aspect ratio $h/d$. Panels (ii)-(iv) present top-view SSIM contours at three morphing stages for each geometry through A-C. The nine morphing stages are corresponding to the nine stages in Figures 5-7, respectively. Despite anisotropies in the Shrinky Dink sheet, glue effects, sharp cut boundaries, and inevitable scanning noise, the SSIM contours show close agreement between simulation and experiment, with most regions appearing in yellow to red.
In Figure 8D, solid markers with error bars indicating standard deviations show the evolution of the aspect ratio $h/d$ over time during thermal actuation. The entire morphing process completes in approximately 15 seconds, dramatically faster than conventional 3D-shape fabrication methods, which often require minutes to hours. Each pattern exhibits a slightly different morphing timeline. The six-arm star on the circular sheet (pattern A) begins deforming earliest at around 5s, and its $h/d$ plateaus after 12s. The two-star pattern on the rectangular sheet (pattern B) begins to morph latest at around 12s. The cross pattern on the square sheet (pattern C) starts at 7s, and ultimately reaches an aspect ratio of $h/d>1.2$. Hollow markers denote the numerical simulations obtained by prescribing thermal strains as described in Sections II.2.2 and II.3. Figure 8E plots the mean SSIM values for the nine SSIM contour plots shown in panels (ii)-(iv). The mean SSIM is simply the arithmetic mean $\overline{\mathrm{SSIM}}=\frac{1}{N_{s}^{3}}\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{s}}\sum_{k=1}^{N_{s}}\mathrm{SSIM}(i,j,k)$. All mean SSIM values are around 0.8, indicating close agreement between the numerical simulations and the scanned experimental shapes.
Figure 8: Validation of numerical simulations against experimental results using SSIM. A-C(i) show side views with the definition of the aspect ratio $h/d$. A-C(ii-iv) present top-view SSIM contours at three morphing stages for each pattern. D shows the evolution of $h/d$ during thermal actuation, where solid markers denote experimental means and hollow markers denote numerical simulations. E summarizes the mean SSIM values for the contours in A-C(ii-iv).
III Conclusion
We introduced a new energy formulation for reduced-order modeling of composites as a single layer material, enabling efficient simulation of stimulus-driven morphing across a broad range of systems beyond bilayers. The formulation couples out-of-plane bending with in-plane stretching, expressed as functions of the applied stimulus, allowing accurate modeling of permanent and stimulus-induced deformations.
We also developed a robust fabrication protocol to transform planar composites into 3D structures via thermal actuation, achieving a repeatable success rate exceeding 90%. Three representative geometries – a six-arm star on a circle, a two-star on a rectangle, and a cross on a square – showed excellent agreement between numerical simulations and experiments across various morphing stages, as evidenced by both qualitative comparisons in Figures 1, 5-7, and quantitative SSIM values in Figure 8.
The close agreement between simulations from our reduced-order model and experiments highlights the predictive power and broad potential for fast, low-cost, and precise design and manufacturing of complex structures through emergent morphogenesis.
IV Experimental Section
Materials and Methods: We used commercially available Shrinky Dink polystyrene sheets (Cridoz Shrinky Art Paper Sheets) and 3D-printed PLA kirigami inspired patterns. The kirigami components were designed using Fusion 360 and SolidWorks CAD software solidworks , utilizing Python scripts to rapidly assign and vary geometry. The designs were fabricated with a fused deposition modelling 3D printer (Bambu Lab 1PS) using PLA filament chosen for its thermal resistance during oven experimentation and its mechanical properties. Adhesion between layers was achieved using a two-part Gorilla epoxy adhesive, selected for its bond strength, thermal stability, cost, and ease of use.
Shrinky Dink Properties: Shrinky Dink is a thermoplastic polystyrene polymer that exhibits isotropic in-plane shrinkage when heated. We define the shrink ratio as $L/L_{0}$, where $L$ is the instantaneous length and $L_{0}$ is the original length. Figure 3 shows the variation of normalized length $\bigl{(}L/L_{0}\bigr{)}$ with normalized temperature $T/T_{g}$. The shrink ratio decreases with increasing temperature and plateaus once $T/T_{g}$ exceeds 1.15, indicating isotropic behavior as the shrink ratios along orthogonal in-plane directions converge. Above $300\,^{\circ}\mathrm{F}$, the $x$- and $y$-axis gauge lengths coincide and the standard deviation bars contract, reflecting highly reproducible shrinkage. To ensure uniform and predictable actuation, all subsequent experiments were conducted within this temperature regime.
Manufacturing Protocol: The fabrication followed four key steps:
  1. 1.
    Print the 3D kirigami design on a fused deposition modeling printer.
  2. 2.
    Bond the 3D printed kirigami elements to pre-cut Shrinky Dink sheets with a thin, uniform layer of epoxy.
  3. 3.
    Cure under distributed weight $(6.8\times 10^{2}\ \mathrm{N/m^{2}})$ for 24 h at room temperature to ensure full adhesion.
  4. 4.
    Thermally activate the composite in a laboratory oven preheated to $300\,^{\circ}\mathrm{F}$. Each sample was placed with the Shrinky Dink side facing down on a non-stick foil base and heated for 15s. The 3D morphing process was observed visually, and samples were removed by hand immediately upon full shape transformation.
Pattern Design and Results: Morphing samples were generated via a Python script that produced random polygonal kirigami patterns on a 2D domain. Visual inspection of the resulting 3D shapes revealed that local curvature propagated along the printed line directions. In a parametric study, we varied three parameters: line width, PLA layer thickness, and line count per polygon. Empirically, increasing line count yielded smoother and more continuous circumferential curvature, whereas fewer lines produced segmented outlines. Thicker and wider PLA lines resisted actuation, requiring longer oven exposure due to higher bending and stretching stiffness. Consequently, designs with more than four lines were adopted for rapid, reliable thermal response and consistent circumferential curvature across samples. Three planar patterns were therefore designed to benchmark the aforementioned observations and to investigate the numerical accuracy of simulations in comparison to the experimentally fabricated patterns. They are: A. Single-star: Fabricated by intersecting three coplanar “arms”, each $90\,\mathrm{mm}$ long, at $60^{\circ}$ increments through the centroid, generating six $45\,\mathrm{mm}$ radial arms. B. Double-star: Superimposed two such stars aligned coaxially, separated by a $56.73\,\mathrm{mm}$ spine, giving an overall tip-to-tip length of $146.7\,\mathrm{mm}$. C. Orthogonal X-cross: Incorporating four arms spaced at $90^{\circ}$ intervals, each extending $56.7\,\mathrm{mm}$ from the center of the shape and spanning $99.99\,\mathrm{mm}$ along the principal diagonal, with each arm chamfered at its end via the addition of a right-angled triangle.
Computational Reconstruction: The computational reconstruction utilized Polycam photogrammetry software alongside a high-resolution camera and a custom-printed stand to capture 80 images around each sample from multiple angles. These images were processed in Polycam to generate and edit 3D models in STL and OBJ formats. To enhance reconstruction consistency and spatial accuracy, we applied fiduciary markers to the samples prior to imaging, facilitating reliable feature recognition and reducing computation time surmen2023 . The resulting meshes were scaled to real-world dimensions based on known physical measurements and subsequently 3D-printed to validate scale fidelity against experimental results.
Aspect Ratio: To extract the mean aspect ratio of each experimental sample, we conducted a frame-by-frame analysis at 1 s intervals using a high resolution video of the thermal actuation procedure. Frames were imported into ImageJ imagej2025 , where a threshold mask isolated the sample from the background for accurate perimeter and geometry detection. The straight-line tool was used to manually define key dimensions, whose pixel lengths were recorded. Conversion from pixels to physical units was performed using a known reference length inside the oven. From these measurements, in-plane shrinkage and out-of-plane height expansion were computed and expressed as the numerical aspect ratio $h/d$, as demonstrated in Figure 8.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author. The data and codes are available on GitHub: https://github.com/StructuresComp/discrete-shells-shrinky-dink
Acknowledgments This work was supported by Samsung Electronics Co., Ltd., and the National Science Foundation under Grants CMMI-2332555, CAREER-2047663, and CMMI-2209782.
Conflict of Interest The authors declare no conflict of interest.
Author Contributions Y.Z. and A.A. contributed equally to this work. Y.Z. developed the numerical methodology. Y.Z., A.A., and M.K.J. designed the experiments. A.A. performed the majority of the experiments and collected most of the data. Y.Z. also contributed to experiments and data collection. Y.Z., A.A., and M.K.J. performed conceptualization, investigation, and formal analysis. M.K.J. acquired funding, and supervised the project. Y.Z., A.A., and M.K.J. wrote, reviewed, and edited the manuscript. All authors read and approved the final manuscript.
Data Availability Statement The data that support the findings of this study are available from the corresponding author upon reasonable request.
V Appendix A. Gradient and Hessian of coupled Energy
We consider axial strain caused by temperature change or prestretch. The thermal strain is given by
$\varepsilon^{\mathrm{th}}=\alpha\,(T-T_{0})$
The gradient and Hessian of the new energy formulation, Equation 6, are given by
$\nabla_{\bm{X}}E(\bm{X})$ $=\sum_{i=0}^{N_{e}-1}k_{s}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{[}\varepsilon_{i}(\bm{X})-\varepsilon_{i}^{\mathrm{th}}\bigr{]}}\,\nabla_{\bm{X}}\varepsilon_{i}(\bm{X})$ (12)
$\quad+\sum_{j=0}^{N_{h}-1}k_{b}\,\bigl{[}\theta_{j}(\bm{X})-\beta l_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{(}\Delta\varepsilon_{j}(\bm{X})-\Delta\varepsilon_{j}^{\mathrm{th}}\bigr{)}}\bigr{]}\,\Bigl{[}\nabla_{\bm{X}}\theta_{j}(\bm{X})-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}(\bm{X})\Bigr{]},$
and
$\nabla_{\bm{X}}^{2}E(\bm{X})$ $=\sum_{i=0}^{N_{e}-1}k_{s}\Bigl{(}\nabla_{\bm{X}}\varepsilon_{i}\otimes\nabla_{\bm{X}}\varepsilon_{i}+\;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{[}\varepsilon_{i}(\bm{X})-\varepsilon_{i}^{\mathrm{th}}\bigr{]}}\,\nabla_{\bm{X}}^{2}\varepsilon_{i}\Bigr{)}$ (13)
$\quad+\sum_{j=0}^{N_{h}-1}k_{b}\Bigl{(}\bigl{[}\nabla_{\bm{X}}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}\bigr{]}\otimes\bigl{[}\nabla_{\bm{X}}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}\bigr{]}$
$\qquad\quad+\,\bigl{[}\theta_{j}(\bm{X})-\beta l_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{(}\Delta\varepsilon_{j}(\bm{X})-\Delta\varepsilon_{j}^{\mathrm{th}}\bigr{)}}\bigr{]}\,\Bigl{(}\nabla_{\bm{X}}^{2}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}^{2}\varepsilon_{jp}\Bigr{)}\Bigr{)}\!.$
Finite-difference approximations have been used to verify the formulas.
VI Appendix B. Dimensions of bilayer composites
The dimensions of the three representative bilayer composites are shown in Figure 9.
Figure 9: Dimensions of the three representative bilayer composites in $m$.